# Trading with Time Series Causal Discovery: An Empirical Study

### v1

@@ Abstract
This study investigates the application of causal discovery algorithms in equity markets, with a focus on their potential to enhance investment strategies. An investment strategy was developed based on the causal structures identified by these algorithms, and its performance was evaluated to assess the profitability and effectiveness in stock market environments. The results indicate that causal discovery algorithms can successfully uncover actionable causal relationships in large markets, leading to profitable investment outcomes. However, the research also identifies a critical challenge: the computational complexity and scalability of these algorithms when dealing with large datasets, which presents practical limitations for their application in real-world market analysis.
@@ 1 Introduction
Numerous causal discovery algorithms have been developed for time series analysis [2], and some have been applied to the stock market to identify underlying driving forces [11]. However, few studies have advanced to the next step of using these driving forces in quantitative studies to predict future stock prices and develop corresponding trading strategies. This report aims to address this gap by designing causality-based trading strategies and evaluating their feasibility and effectiveness.
The major contributions of this report include the following:
@@
    Developing a workflow to translate causality into trading profits.
@@
    Applying this workflow to major stock markets in China and the US.
@@
    Providing clear answers to the client’s three questions.
The Python and R code developed in this research, including data pre-processing scripts, accuracy evaluation tools, causal discovery interfaces, causal graph reformatting tool, auto-regression fitting and prediction routines, trading simulators, backtesting utilities, trading performance visualizers, and execution time profiling tools, is open-sourced [^1] GitHub repository: https://github.com/kaaaylaaa/causality-based-trading.
The remainder of this report is organized as follows: Section 2 provides a brief introduction to the selected causal discovery algorithms, namely tsFCI, VarLiNGAM, and TiMINo. Section 3 outlines the detailed procedures of our workflow. Section 4 presents the experimental results along with their analysis. Finally, Section 5 offers conclusions and recommendations for future research, and answers to the client’s questions.
@@ 2 Background
tsFCI [5] Constraint-Based Causal graph formatting
VarLiNGAM [13] Noise-Based Causal graph formatting
TiMINo [18] Noise-Based Edge direction decoder
@@ 3.1 Data Preparation
We performed data imputation due to the substantial amount of missing values in the datasets. The imputation process involved two steps. First, linear interpolation was applied across all stocks to estimate and fill gaps where missing values occurred between known values. Second, any stock that still contained missing values after interpolation was removed from the dataset. As a result, the cleaned SP500 dataset comprises 446 stocks, while the cleaned CSI300 dataset comprises 98 stocks. The Pelosi dataset contains no missing values, so no stocks were removed. An overview of our data is provided in Table 2.
@@ 3.2 Causal Discovery, Prediction and Trading
where $\tau$ is the time lag used during causal discovery, and we start from $t-1$ instead of $t$. In other words, the current value is modelled by the changes in the driving forces at previous time steps, while the present does not exert an instantaneous effect. An expanding window is employed when fitting the predictive model. The daily prediction and trading process is outlined in Figure 2. For each stock, we extract the historical values of its driving forces, fit a linear regression model on these past prices, and then make a one-step ahead prediction. Using the predicted prices, we calculate the predicted returns for the next day as follows:
@@
Given price predictions, numerous studies have demonstrated that a long-short momentum strategy can generate significant profits [14, 23, 16, 17]. This strategy operates on the premise that stocks with strong past performance are likely to continue performing well, while those with poor past performance are expected to underperform. The cross-sectional long-short momentum strategy introduced by [14] constructs a zero-cost market-neutral portfolio by buying stocks with high past returns (winners) and selling those with low past returns (losers). Building on this approach and integrating short-term forecasting, our trading strategy creates a dollar-neutral portfolio, ensuring that the dollar amounts of long and short positions are balanced [4, 1]. For simplicity, we assume a liquid market where short selling is permitted without incurring borrowing costs. However, given that our strategy operates on a daily basis, transaction costs cannot be ignored as they can significantly impact portfolio returns. Therefore, we impose a fixed daily transaction cost of 0.1% [7]. The strategy is implemented as follows:
@@ 4.2 Discussion
Our results indicate that incorporating a causal discovery phase generally improves portfolio returns, especially with larger datasets, where profitability increases with data size. Among the three algorithms, VarLiNGAM is the most effective, handling large datasets well and generating the highest returns on all datasets. Major observations we obtain from the results include the following:
@@ 5 Conclusion
We finish this report by answering the questions specified by our client:

### v2

@@ Abstract
This study investigates the application of causal discovery algorithms in equity markets, with a focus on their potential to build investment strategies. An investment strategy was developed based on the causal structures identified by these algorithms. The performance of the strategy is evaluated based on the profitability and effectiveness in stock markets. The results indicate that causal discovery algorithms can successfully uncover actionable causal relationships in large markets, leading to profitable investment outcomes. However, the research also identifies a critical challenge: the computational complexity and scalability of these algorithms when dealing with large datasets. This challenge presents practical limitations for their application in real-world market analysis.
@@ 1 Introduction
Numerous causal discovery algorithms have been developed for time series analysis [2], and some have been applied to the stock market to identify underlying driving forces [11]. However, few studies have advanced to the next step of using these driving forces in quantitative studies to predict future stock prices and develop corresponding trading strategies. This paper aims to narrow this gap by designing causality-based trading strategies and evaluating their feasibility and effectiveness.
The major contributions of this paper include the following:
@@
    Applying time series causal discovery to real-world stock data
@@
    Developing a workflow to turn causal relations into a trading strategy
@@
    Backtesting the strategy on major stock markets in China and the US.
The Python and R code developed in this research, including data pre-processing scripts, accuracy evaluation tools, causal discovery interfaces, a causal graph reformatting tool, auto-regression fitting and prediction routines, trading simulators, backtesting utilities, trading performance visualizers, and execution time profiling tools, is open-sourced [^1] GitHub repository: https://github.com/kaaaylaaa/causality-based-trading.
In this study, we mainly focus on three key questions:
  • Q1
    Can we quantitatively evaluate the effectiveness of a causal discovery algorithm using stock price data?
  • Q2
    If effectiveness can be quantified, which causal discovery algorithm performs best in driving force analysis for the market?
  • Q3
    Are there practical challenges when applying causal discovery to analyze stock markets, even if the methods are theoretically sound?
The remainder of this paper is organized as follows: Section 2 provides a brief introduction to the selected causal discovery algorithms, namely tsFCI, VarLiNGAM, and TiMINo. Section 3 outlines the detailed procedures of our workflow. Section 4 presents the experimental results along with their analysis. Finally, Section 5 offers conclusions and recommendations for future research, and answers to the research questions.
@@ 2 Background
tsFCI [5] Constraint-Based Causal graph reformatting
VarLiNGAM [13] Noise-Based Causal graph reformatting
TiMINo [18] Noise-Based Edge direction determination
@@ 3.1 Data Preparation
We performed data imputation due to the substantial amount of missing values in the datasets. The imputation process involved two steps. First, linear interpolation was applied across all stocks to estimate and fill gaps where missing values occurred between known values. Second, any stock that still contained missing values after interpolation was removed from the dataset. As a result, the cleaned SP500 dataset comprises 446 stocks, while the cleaned CSI300 dataset comprises 98 stocks. The Pelosi dataset contains no missing values, so no stocks were removed. An overview of our data is in Table 2.
@@ 3.2 Causal Discovery, Prediction and Trading
where $\tau$ is the time lag used during causal discovery. The latest time point we use for prediction is $t-1$ instead of $t$. In other words, the current value is modelled by the changes in the driving forces at previous time steps, while the present does not produce an instantaneous effect. An expanding window is employed when fitting the predictive model. The daily prediction and trading process is outlined in Figure 2. For each stock, we extract the historical values of its driving forces, fit a linear regression model on these past prices, and then make a one-step ahead prediction. Using the predicted prices, we calculate the predicted returns for the next day as follows:
@@
Given price predictions, numerous studies have demonstrated that a long-short momentum strategy can generate significant profits [14, 23, 16, 17]. This strategy operates on the premise that stocks with strong past performance are likely to continue performing well, while those with poor past performance are expected to underperform. The cross-sectional long-short momentum strategy introduced by [14] constructs a zero-cost market-neutral portfolio by buying stocks with high past returns (winners) and selling those with low past returns (losers).
Based on this approach and integrating short-term forecasting, our trading strategy employs a dollar-neutral portfolio, ensuring that the dollar amounts of long and short positions are balanced [4, 1]. For simplicity, we assume a liquid market where short selling is permitted without incurring borrowing costs. However, given that our strategy operates on a daily basis, transaction costs cannot be ignored as they can significantly impact portfolio returns. Therefore, we impose a fixed daily transaction cost of 0.1% [7]. The strategy is implemented as follows:
@@ 4.2 Discussion
Our results indicate that incorporating a causal discovery phase generally improves portfolio returns, especially with larger datasets. Among the three algorithms, VarLiNGAM is the most effective, handling large datasets well and generating the highest returns on all datasets. Major observations we obtain from the results include the following:
@@ 5 Conclusion
We finish this paper by answering the research questions specified in Section 1: